OpenStudy (anonymous):

Samantha has 86 cents. She gives 31 cents to her friend, then accidentally loses a fifth of what is left. What fraction of the original 86 cents does she still have?

OpenStudy (anonymous):

I came up with 44 cents.

OpenStudy (anonymous):

i got that too but i need it in a fraction

OpenStudy (anonymous):

11/25 I think.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yeah, no problem

OpenStudy (anonymous):

\[\frac{86-31-\left(\frac{86-31}{5}\right)}{86}=\frac{22}{43} \]

OpenStudy (anonymous):

@Gin_Ichimaru Thank you for the medal.

OpenStudy (anonymous):

No Prob @robtobey